Can you please give me a step by step on how to do the following 5 questions...
1.You and your cousin go to Wendy's for a "big" lunch. You buy 3 burgers and 2 orders of fries for $6.50. Your cousin buys 2 burgers and 5 orders of fries for $8.00. How much did each item cost?
2.Jermaine has a jar of nickles and dimes. There are 200 coins worth $14.00. How many of each type of coins are in the jar?
3.The perimeter of a rectangle garden is 96m. The legnth of the garden is 8 more than the width. Find the legnth and width of the garden.
3.The radio station has 55 employees. There is 1 less than three times as many men as women working at the station. How amny indiviual men and women work at the station?
4.During the hoidays, your family buys boxes of candy canes and chocolate covered cherries. You bought 25 boxes of candy and spent $28.00. If the cherries cost $1.00 a box and the candy canes cost $1.30, how many of each type of candy did you buy?
5.The sumof two numbers is 8. One number is 3 more than 4 times the other. What are the numbers?
Thank You for your time please email me back as soon as possible!!!

Answer by stanbon(75887)

1.You and your cousin go to Wendy's for a "big" lunch. You buy 3 burgers and 2 orders of fries for $6.50. Your cousin buys 2 burgers and 5 orders of fries for $8.00. How much did each item cost?
-----
Equations:
3b+2f = 650 cents
2b+5f = 800 cents
------
Multiply thru 1st by 2
Multiply thru 2nd by 3
---
6b + 4f = 1300
6b +15f = 2400
-----
Subtract 1st from 2nd and solve for "f":
11f = 1100
f = 100 cents (The price of fries is $1.00)
----
Solve for "b":
2b + 5f = 800
2b + 5*100 = 800
2b = 300
b = 150 cents (The price of a burger is $1.50)
=======================
2.Jermaine has a jar of nickles and dimes. There are 200 coins worth $14.00. How many of each type of coins are in the jar?
---
Quantity Eq: n + d = 200
Value Eq: 5n+10d = 1400 cents
------------------
Multiply thru the Quantity Eq. by 5:
5n + 5d = 1000
------
Subtract that from the Value Equation: and solve for "d":
5d = 400
d = 80 (number of dimes)
---
Solve for "n":
n + d = 200
n + 80 = 200
n = 140 (number of nickels)
================================
The following problems also have simple equations that
can be solved by elimination.
Please give it a try.
Cheers,
Stan H.
================
3.The perimeter of a rectangle garden is 96m. The length of the garden is 8 more than the width. Find the length and width of the garden.
3.The radio station has 55 employees. There is 1 less than three times as many men as women working at the station. How amny indiviual men and women work at the station?
4.During the hoidays, your family buys boxes of candy canes and chocolate covered cherries. You bought 25 boxes of candy and spent $28.00. If the cherries cost $1.00 a box and the candy canes cost $1.30, how many of each type of candy did you buy?
5.The sumof two numbers is 8. One number is 3 more than 4 times the other. What are the numbers?